### v2

# astroCAMP: A Community Benchmark and Co-Design Framework for Sustainable SKA-Scale Radio Imaging

Abstract.
The Square Kilometre Array (SKA) will operate one of the world’s largest continuous scientific data systems, sustaining petascale imaging under strict power envelopes. Yet current radio-interferometric pipelines typically achieve only 4–14% of hardware peak because of memory and I/O bottlenecks, resulting in high energy, operational, and carbon costs. Progress is further constrained by the absence of standardised cross-layer metrics and survey-level fidelity tolerances for principled hardware–software co-design. We present astroCAMP, a reproducible benchmarking and co-design framework for SKA-scale imaging. astroCAMP contributes: (1) a unified metric suite spanning performance, utilisation, memory/data-movement behavior, sustainability, economics, and scientific fidelity; (2) standardised SKA-representative datasets, reference outputs, and benchmark configurations for reproducible cross-platform evaluation; (3) a multi-objective co-design formulation linking quality constraints to time-, energy-, carbon-, and cost-to-solution; and (4) a reproducible design-space exploration workflow to derive Pareto-optimal operating regions. We release datasets, scripts, benchmark results, and a reproducibility kit, and evaluate WSClean+IDG on an AMD EPYC 9334 CPU and an NVIDIA H100 GPU. The evaluation shows substantial end-to-end orchestration and synchronization bottlenecks despite efficient kernels in active phases, limited CPU strong scaling, and location-dependent carbon/cost efficiency under realistic grid and electricity-price assumptions. We further illustrate the use of astroCAMP for heterogeneous CPU–FPGA design-space exploration, and its potential to facilitate the identification of Pareto-optimal operating points for SKA-scale imaging deployments. Lastly, we call on the SKA community to define quantifiable fidelity metrics and thresholds to accelerate principled optimisation for SKA-scale imaging.
radio astronomy, interferometric imaging, benchmarking, co-design, HPC, energy efficiency, carbon-aware computing, reproducibility
copyright: noneconference: Conference; Switzerland;
1. Introduction
Modern radio interferometers are entering a regime where computing, not photon collection, limits scientific capability. Arrays such as LOFAR, MWA, ASKAP, and the VLA already generate visibility, calibration, and imaging workloads that strain current HPC facilities, and these pressures intensify for the Square Kilometre Array (SKA), the most data-intensive radio observatory ever built. SKA-Low and SKA-Mid—illustrated in Fig. 1—will stream $\sim$8–20 Tb/s of correlated visibilities to two Science Data Processors (SDPs), each expected to operate within a site-level power cap of 5 MW for several decades (Dewdney et al., 2009; Broekema et al., 2015). Because both SDPs are located on carbon-intensive grids in South Africa and Australia, energy efficiency directly determines operational cost, scientific throughput, and long-term sustainability.
Figure 1. SKA’s infrastructure includes 2 Central Signal Processors (CSPs), 2 Science Data Processors (SDPs), and a global network of data centers (SRCNet). The SKA-Low and SKA-Mid telescopes will stream 28 Tb/s of data to the CSPs, and the SDPs will process it into science data products.
Efficiency requirements for SKA SDPs. Public SKA design documents (Dewdney et al., 2009; Broekema et al., 2015) indicate that the Science Data Processor (SDP) must sustain approximately 16–42 PFLOP/s for SKA-Low and 20–72 PFLOP/s for SKA-Mid, depending on observing mode [^1] Online SKA SDP performance-model references: https://developer.skao.int/projects/sdp-par-model/en/latest/ and https://ska-telescope.gitlab.io/sdp/ska-sdp-par-model/notebooks/SKA1_Imaging_Performance_Model.html.. For a sustained compute requirement $C_{\rm sus}$ and power envelope $P_{\rm cap}$, the facility-level efficiency is $\eta_{\rm facility}=\frac{C_{\rm sus}}{P_{\rm cap}}.$ Only a fraction of this power is available to compute nodes, as modern HPC systems incur 30–40% overheads for networking, memory, storage, and cooling (PUE 1.3–1.4). Adopting $f_{\rm overhead}\in[0.3,0.4]$, the compute-node requirement becomes $\eta_{\rm compute}=\frac{C_{\rm sus}}{P_{\rm cap}(1-f_{\rm overhead})}.$
Fig. 2 summarizes the required efficiency ranges for SKA-Low and SKA-Mid under 1–5 MW power caps. A 2 MW allocation for SKA-Low and SKA-Mid requires tens of GFLOP/s/W, comparable to the average Green500 systems in November 2025. These typically deliver 30–40 GFLOP/s/W. KAIROS, the current Green500 leader, reaches 73.28 GFLOP/s/W (TOP500.org, 2025). This is comparable to the most computationally demanding operation case for SKA-Low, under a power cap of 1 MW. However, Green500 benchmarking numbers reflect idealised, compute-bound kernels (LINPACK), whereas SKA imaging workloads (non uniform 2D and 3D FFTs) are in theory memory-limited and in practical implementations I/O-limited. Prior works in major algorithmic advances—such as $w$-projection (Cornwell et al., 2008), $w$-stacking (Pratley et al., 2019), faceting, and IDG (Van der Tol et al., 2018b)—radio imaging remain dominated by memory bandwidth (FFTs), data movement, and irregular access patterns.
The utilisation gap for imaging software deployments. On modern CPU and GPUs, imaging pipelines typically sustain only 4–14 % of peak floating-point performance (Veenboer and Romein, 2020), revealing a fundamental mismatch between imaging workloads and commodity architectures, as also indicated by the pink band in Fig. 2. Earlier modelling (Broekema et al., 2015) reported only 10–20% efficiency for FFT- and gridding-dominated workloads, implying that naïvely scaling current pipelines to SKA data rates would require tens of megawatts, well beyond SDP site caps. Without major improvements in software efficiency and hardware utilisation, meeting SKA throughput would require substantial overprovisioning and likely exceed the power budget. The core issue is that inefficient implementations can draw significant power even when compute units are under-utilised, producing high operational and carbon costs for limited scientific return.
[…]
Closing this gap requires more than incremental optimization of hardware and software tuning. It demands holistic co-design across algorithms, data structures, and hardware:
  1. (1)
    Algorithmic optimisation: reduce data movement through hierarchical memory reuse, structured sparsity, and compressed visibility formats.
  2. (2)
    Domain-specific accelerators: exploit GPUs, FPGAs, custom architectures, and ASICs optimised for imaging hotspots.
  3. (3)
    Energy-aware orchestration: use dynamic power scaling and locality-aware scheduling to match data-access patterns.
Principled co-design further requires standardised datasets, cross-layer metrics, science-driven image-quality tolerances, and reproducible evaluation methods under representative SKA workloads. Without such framework, optimisation risks becoming a blind pursuit of efficiency gains, obscuring the central question for SKA-scale computing: how much scientific fidelity can be traded for performance, energy, and carbon savings without compromising discovery?
Machine learning, for example, overcame a similar lack of standardisation through community benchmarks such as MLPerf (Reddi et al., 2020; Farrell et al., 2021), unified datasets, accuracy thresholds, and submission protocols. In large-scale computational physics, mixed-precision solvers exploit application-level convergence tolerances to improve performance while preserving required scientific accuracy (Clark et al., 2010). In climate modeling, the Coupled Model Intercomparison Project (CMIP) defines standardized experimental protocols and evaluation criteria to ensure comparability across models (Eyring et al., 2016). Radio astronomy lacks an equivalent effort. These gaps hinder agile innovation in the domain and prevent researchers and developers from comparing figures of merit without re-running large, expensive experiments.
Scope and contributions of this work. Towards bridging these gaps, this paper introduces astroCAMP, a reproducible benchmarking and co-design framework for radio-interferometric imaging. astroCAMP provides the metrics, datasets, and benchmark cases required to evaluate efficiency, sustainability, and quality trade-offs across high-performance and heterogeneous architectures. The contributions are:
  1. (1)
    A unified cross-layer metric suite. We define and implement a set of 12+ measurable metrics spanning performance, energy, carbon, system behaviour, economic cost, and scientific fidelity, providing a consistent basis for comparing imaging pipelines across CPUs, GPUs, FPGAs, ASICs, and emerging accelerators.
  2. (2)
    Standardised SKA benchmark suite and datasets. We provide visibility datasets, reference dirty images, and parameterised benchmark configurations in the astroCAMP GitHub repository (SEAMS Project, 2025). With this, we aim to enable reproducible, cross-platform evaluation of SKA imaging pipelines and systematic exploration of hardware–algorithm co-design trade-offs.
  3. (3)
    A multi-objective co-design formulation. We formalise imaging as an optimisation problem over algorithmic and architectural parameters, with the objective to minimize time- energy- and carbon-to-solution under explicit quality, cost, and power constraints.
  4. (4)
    A reproducible design-space exploration workflow. Using the PREESM framework (Suriano et al., 2017), we demonstrate for a subset of astroCAMP’s metrics suite how astroCAMP can enable multi-objective design-space exploration and the computation of Pareto fronts in the metrics domain space.
Together, the four contributions form the foundation of the first end-to-end methodology for rigorously evaluating and co-designing SKA-scale imaging pipelines with carbon efficiency as first-class objective.
The remainder of this paper is organized as follows. Section 2 outlines the sustainability goals and challenges of the SKA SDP. Section 3 presents the astroCAMP framework and co-design methodology. Section 4 introduces the multi-objective co-design formulation and Section 5 the benchmark suite. Section 6 reports the experimental evaluation and results, while Section 7 discusses their implications for hardware–software co-design. Section 8 concludes the paper.
2. SKA SDP Sustainability Challenges and Gaps
Why does optimizing efficiency matter? Recent studies highlight the ecological impact of large-scale scientific computing in astronomy (Portegies Zwart, 2020; Aujoux et al., 2021; dos Santos Ilha et al., 2024; Knödlseder et al., 2022). In response, the SKA Observatory has embedded sustainability and net-zero objectives into its 50-year roadmap, aligned with the UN Sustainable Development Goals, and adopted the CO2 Performance Ladder to support its net-zero transition (SKAO Communications, 2024b, a).
[…]
Because operational emissions scale directly with compute efficiency, the limitations of current platforms become critical. In practice, for SKA-scale workloads, low arithmetic intensity and utilisation mean:
  • Scalability failures: many imaging workloads encounter strong-scaling limits in practice due to communication, synchronization, and serial overheads, yielding diminishing throughput gains as hardware is added.
  • Higher operational and capital costs: more nodes are required to offset low utilisation, increasing electricity and cooling demands.
  • Higher carbon emissions: fewer images, catalogues, and time series are delivered per ton of CO2 emitted.
Energy efficiency, utilisation, and parallel scaling are therefore not just performance metrics but climate performance indicators, coupling algorithmic design and hardware co-design to measurable carbon reduction. A factor-of-two improvement in application-level GFLOP/s/W directly halves SDP operational emissions at fixed science throughput. Moreover, utilization improvement reduces total cost of ownership by lowering both CAPEX and OPEX (Rodríguez Álvarez et al., 2025).
2.1. Gaps in Current Evaluation Practice
Table 1. Metric coverage across imaging tools: pipelines (WSClean, DDFacet, BIPP, WS-Snapshot), algorithmic kernels (IDG), and hardware kernels (FPGA-RP). Symbols: explicit (✓), implicit ($\sim$), missing (—). Metrics are detailed in Table 2.
Metric Category WSClean IDG DDFacet FPGA BIPP WS-S. A.CAMP
Pub. Year 2014 2018 2020/23 2022 2025 2025
System-level (heterogeneous node / pipeline)
Time-to-solution $T_{c}$
Energy-to-solution $E_{c}$ $\sim$
Throughput $\Theta$ (vis/s) $\sim$ $\sim$
Energy efficiency $\eta_{E}$ $\sim$
Hardware platform-level (CPU / GPU / FPGA / ASIC)
Utilisation / occupancy $U$ $\sim$
Memory bandwidth $B_{\mathrm{mem}}$
Memory efficiency $\eta_{mem}$ $\sim$
Peak memory usage $M_{\mathrm{peak}}$ $\sim$ $\sim$ $\sim$ $\sim$ $\sim$
Sustainability
Carbon-to-solution $C_{c}$
Carbon efficiency $\eta_{C}$
Economics
Cost per job
Recent wide-field imaging studies have delivered significant advances in algorithmic sophistication and numerical fidelity, yet their evaluation methodology remains largely performance-centric. As summarised in Table 1, runtime and dirty-image RMS are consistently reported across WSClean (Offringa et al., 2014), BIPP (Tolley et al., 2025), WS-Snapshot (Wu et al., 2025), IDG (Van der Tol et al., 2018b), FPGA-RP (Corda et al., 2022b), and DDFacet (Tasse et al., 2023; Monnier et al., 2020), but key architecture-level metrics—power, energy-to-solution, energy efficiency, roofline characterisation, bandwidth sensitivity, and hardware utilisation— are either missing or only implicitly analysed. Even in DDFacet, where multi-node parallelisation has been demonstrated (Monnier et al., 2020), no measurements of power or energy per job are reported. This limits and slows quantitative comparison across CPUs, GPUs, FPGAs, and emerging accelerators towards optimizing efficiency metrics.
System-level behaviour exhibits similar fragmentation. Pipeline-stage breakdowns and parallel efficiency are available in DDFacet’s distributed implementation (Monnier et al., 2020), and partially in IDG (Van der Tol et al., 2018b) and FPGA-RP (Corda et al., 2022b), but metrics such as device occupancy, data locality, bandwidth limits, and end-to-end dataflow behaviour are rarely treated systematically. This hinders the identification of system bottlenecks that directly influence scalability and energy efficiency in SKA-scale deployments.
None of the surveyed works reports carbon-to-solution, carbon efficiency, or any economic metrics, despite their increasing relevance given SKA’s strict power and environmental constraints. Overall, the literature reveals a fragmented evaluation landscape in which no existing study jointly assesses architecture-level efficiency, system-level behaviour, algorithmic fidelity, and sustainability. To date, only one imaging study has reported total energy-to-solution (Corda et al., 2022b), but only at the algorithmic kernel level (IDG), not for the entire imaging pipeline (WSClean). astroCAMP addresses this gap by providing a unified, reproducible, and cross-layer metric framework that enables principled algorithm–hardware co-design and supports energy-, carbon-, and fidelity-aware optimisation for large-scale imaging systems.
3. astroCAMP Framework
astroCAMP provides the missing infrastructure for reproducible, cross-layer evaluation of radio-interferometric imaging software and hardware. It unifies algorithmic, architectural, and sustainability metrics into a single benchmarking and co-design workflow (Fig. 3), enabling transparent comparison of imaging algorithms and hardware platforms. The framework enables designers to examine:
  1. (1)
    how algorithmic parameters (precision, kernel sizes, tiling) interact with CPU/GPU/FPGA platforms, supporting Outcome 1 by producing reusable, cross-platform performance and fidelity baselines;
  2. (2)
    which configurations satisfy SKA SDP power and throughput limits and populate Pareto-optimal performance–energy–quality trade spaces, delivering Outcome 2; and
  3. (3)
    how close commodity hardware can approach SKA efficiency targets and where accelerators are justified, directly informing Outcome 3 on energy- and carbon-efficient co-design.
Figure 3. astroCAMP co-design framework. Dotted modules are intended for community extension.
3.1. Benchmarking and Co-Design Methodology
astroCAMP provides a reproducible methodology for evaluating radio-interferometric imaging pipelines across heterogeneous architectures. The framework standardises benchmark cases, input–output configurations, and a cross-layer metric suite (performance, energy, carbon, utilisation, and scientific fidelity), enabling fair comparison of algorithmic variants and hardware platforms under realistic SKA-class constraints. The proposed methodology ensures that results obtained on different systems and with different imaging algorithms remain directly comparable. It consists of three non-overlapping stages (illustrated in Fig. 3):
1. Profiling & Measurements. Each benchmark is executed under controlled and repeatable conditions. We wait for node to reach a steady thermal and power state. Rack-level power distribution units (PDUs) provide energy measurements to calculate energy-to-solution $E_{c}$, while platform telemetry (e.g., PMT (Corda et al., 2022a) for CPUs, NVML for GPUs) is timestamp-aligned with PDU traces to validate device-level behaviour. From these measurements we derive time-to-solution $T_{c}$, throughput, and carbon-to-solution $C_{c}$ using region-specific carbon intensity. All scripts, configurations, and reproducibility guidelines are openly available (SEAMS Project, 2025).
2. Co-Design Metrics Evaluation. For each benchmark configuration—including algorithmic parameters (e.g., IDG kernel sizes) and platform choice (CPU, GPU, FPGA)—we assemble a consistent vector of cross-layer metrics defined in Section 3.2. These metrics capture performance, energy, cost, sustainability, and system constraints (e.g., SKA SDP node power caps). This stage performs metric computation only: it normalises and validates measurements across systems, preparing a unified metric vector that serves as input to optimisation. In the current release, system and platform metrics are evaluated using PREESM (Suriano et al., 2017), and sustainability/economic metrics via CEO-DC (Rodríguez Álvarez et al., 2025). Algorithm-level quality metrics are intentionally excluded in this release and form an open call for the community to converge on standardised tolerances.
3. Co-Design Loop & Multi-Objective Optimisation. The metric vectors from Stage 2 feed into a structured design-space exploration (DSE) loop following the formulation in Section 4. The objective is to systematically evaluate alternative algorithm–architecture mappings and solve the resulting multi-objective optimisation problem to produce reproducible Pareto frontiers and trade-off curves. This stage provides the interpretation layer: identifying efficient regions of the design space, exposing performance–energyquality bottlenecks, and determining when approximate methods or domain-specific accelerators become necessary to meet SKA-scale efficiency targets.
3.2. Co-design Metrics
Table 2. astroCAMP co-design metrics. Core metrics (C) enter the performance–energy–quality-cost optimisation; diagnostic metrics (D) support interpretation and reproducibility. Type symbols: Direct = $\bullet$, Derived = $\circ$, Proxy/Model-based = $\triangle$.
Layer Metric / Formula Unit Role Type Instrumentation Interpretation
System (Pipeline) Time-to-solution $T_{c}$ s C $\bullet$ POSIX time, workflow logs, scheduler timestamps End-to-end wall-clock runtime of the workload on the full (possibly heterogeneous) system, where $T_{c}$ is the elapsed execution time.
Energy-to-solution $E_{c}=\int_{0}^{T_{c}}P(t)\,dt$ J C $\bullet$ Rack/node PDUs, CPU RAPL, GPU NVML, PMT traces Total electrical energy consumed by the system, integrating instantaneous power $P(t)$ over runtime $T_{c}$.
Throughput $\Theta=N/T_{c}$ vis/s C $\circ$ MS/FITS logs, visibility/image counters System-level processing rate, where $N$ is the number of visibilities or images processed during runtime $T_{c}$.
Energy efficiency $\eta_{E}=N/E_{c}$ vis/J C $\circ$ Derived from $N$ and $E_{c}$ System-level science throughput per joule, using processed data volume $N$ and energy-to-solution $E_{c}$.
Data locality $L_{d}=V_{\mathrm{local}}/V_{\mathrm{total}}$ D $\triangle$ Darshan I/O profiler, filesystem telemetry Fraction of total I/O volume served locally, where $V_{\mathrm{local}}$ and $V_{\mathrm{total}}$ are local and total I/O volumes.
Hardware Platform Utilisation $U=t_{\mathrm{active}}/t_{\mathrm{total}}$ D $\bullet$ nvidia-smi, ROCm tools, perf, Prometheus exporters Fraction of time a device (CPU socket, GPU, FPGA, etc.) is actively running kernels, where $t_{\mathrm{active}}$ is active compute time and $t_{\mathrm{total}}$ is total wall time.
Memory bandwidth $B_{\mathrm{mem}}=\mathrm{Bytes}/T_{c}$ GB/s D $\bullet$ Hardware counters, Intel VTune, NVIDIA Nsight Sustained data-movement rate per platform, with $B_{\mathrm{mem}}$ indicating whether kernels are memory-bound.
Memory efficiency $\eta_{\mathrm{mem}}=\mathrm{Bytes}/E_{c}$ GB/J D $\bullet$ Derived from memory-traffic counters and energy-to-solution $E_{c}$ Useful data movement per joule at system level, computed as total DRAM bytes transferred divided by system energy-to-solution $E_{c}$. Complements $B_{\mathrm{mem}}$ by quantifying the energy cost of memory traffic.
Peak memory usage $M_{\mathrm{peak}}$ GB D $\bullet$ /proc/meminfo, nvidia-smi, cgroups, container telemetry Maximum resident memory footprint observed on a platform during execution, constraining batch size and problem scaling.
Algorithmic Quality Dirty-image RMS $\sigma_{\mathrm{dirty}}=\sqrt{\frac{1}{N}\sum(I_{i}-\bar{I})^{2}}$ Jy/beam C $\bullet$ CASA/WSClean imstat, PyBDSF Noise and artefact level in the dirty image, using pixel intensities $I_{i}$, mean value $\bar{I}$, and resulting RMS $\sigma_{\mathrm{dirty}}$.
PSNR / SSIM, PSNR $10\log_{10}(I_{\max}/\mathrm{MSE})$ dB / – C $\bullet$ scikit-image, OpenCV Fidelity of reconstruction $\hat{I}$ vs. reference $I_{\mathrm{ref}}$, using maximum pixel $I_{\max}$ and mean-squared error; SSIM measures structural similarity between $\hat{I}$ and $I_{\mathrm{ref}}$.
Dynamic range $DR=I_{\max}/\sigma_{\mathrm{res}}$ C $\bullet$ Residual-image statistics Ratio of peak intensity $I_{\max}$ to residual RMS $\sigma_{\mathrm{res}}$, determining detectability of faint emission.
Astrometric error $\epsilon_{\mathrm{astro}}=\frac{1}{N}\sum\|\mathbf{x}_{i}-\mathbf{x}_{i}^{\mathrm{ref}}\|$ arcsec or px C $\bullet$ PyBDSF catalogues Average positional discrepancy between reconstructed and reference sources, where $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{\mathrm{ref}}$ are measured and reference positions.
Sustain. Carbon-to-solution $C_{c}=E_{c}\,\kappa(t,r)$ gCO2e C $\triangle$ Grid carbon-intensity APIs (e.g. electricityMap) + $E_{c}$ Total carbon footprint of a run, using measured energy $E_{c}$ and grid carbon intensity $\kappa(t,r)$ at time $t$ and region $r$.
Carbon efficiency $\eta_{C}=N/C_{c}$ vis/gCO2e D $\circ$ Derived from $N$ and $C_{c}$ Visibilities processed per gram CO2 emitted, normalising science throughput $N$ by carbon cost $C_{c}$.
Economic Total cost ownership $C_{\mathrm{TCO}}=C_{\mathrm{capex}}+C_{\mathrm{opex}}$ D $\bullet$ Procurement & operations records Lifetime system cost, combining capital cost $C_{\mathrm{capex}}$ and operational cost $C_{\mathrm{opex}}$.
Cost per job $C_{E}=E_{c}\,p_{E}$ C $\circ$ Electricity tariff + $E_{c}$ Monetary cost per workload, using energy-to-solution $E_{c}$ and electricity price $p_{E}$.
Cost efficiency $\Theta/C_{\mathrm{TCO}}$ vis/€ D $\circ$ Throughput + TTO data Science operations delivered per euro invested, using throughput $\Theta$ and total cost $C_{\mathrm{TCO}}$.
astroCAMP includes a compact, cross-layer metric suite that links hardware behaviour, energy use, economic cost, scientific throughput, and fidelity. Table 2 summarises all metrics, including their formulas, units, role, and instrumentation, and interpretation. The objective function is to minimize time-, energy-, carbon, and cost-to-solution while maximizing science throughput. Only the core metrics (C) are target for optimization, while diagnostic metrics (D) support the interpretation of results by revealing system-level and platform-level bottlenecks—such as remote data access, uneven work distribution across CPUs or GPUs, and delays caused by slower devices.
The framework captures system-level (pipeline) key performance indicators, including time-to-solution $T_{c}$, energy-to-solution $E_{c}$, and scientific throughput (visibilities/s or images/s), as well as hardware platform (architecture) metrics to inform co-design decisions, such memory bandwidth, memory efficiency, and utilisation. Algorithmic quality is quantified via dirty-image RMS, PSNR/SSIM, dynamic range, and astrometric and photometric errors from catalogue comparisons, capturing the scientific impact of approximations such as reduced precision, coarse $w$-stacking, or NUFFT kernel truncation. Sustainability metrics pair energy use with regional carbon intensity $\kappa(t,r)$ to compute carbon-to-solution $C_{c}$, enabling joint optimisation of energy, carbon, and image fidelity. Economic metrics—including cost per job $C_{E}=E_{c}\,p_{E}$ and amortised total cost of ownership $C_{\mathrm{TCO}}$—connect computational choices to financial constraints relevant for large-scale SKA deployments. Together, the full metric suite supports transparent, multi-objective co-design across CPU, GPU, FPGA, and ASIC platforms.
4. Extensible Multi-Objective Co-Design Formulation
astroCAMP provides a general and extensible co-design formulation that unifies algorithmic, architectural, and system-level degrees of freedom. Rather than prescribing specific algorithms or hardware, astroCAMP defines a stable metric backbone (Table 2) onto which current and future imaging algorithms (e.g., WSClean, IDG, BIPP), accelerator technologies (CPUs, GPUs, FPGAs, ASICs), and workflow runtimes can be instantiated.
Design space. Any imaging configuration is represented as a design point $x=(a,h,s)\in\mathcal{X}$, shown in Fig. 4, where:
  • $a$ are algorithmic parameters: precision, convolution kernel size, NUFFT or $w$-stacking order, snapshot parameters, visibility tiling, task fusion.
  • $h$ are hardware parameters: processor type, accelerator count, memory hierarchy, power caps, thread/block geometry.
  • $s$ are system/workflow parameters for mapping $a$ to $h$: parallel decomposition, data-locality strategy, I/O staging, buffering and batching, workflow orchestration.
Design Space x=(a,h,s)x=(a,h,s)Algorithmic aaHardware hhSystem ssAll feasible configurationsMetric evaluationObjective Space $(T_{c},E_{c},C_{c},C_{TCO},\Delta Q)$$T_{c}$$E_{c}$Pareto frontier(Projection of 5D objective space)
Figure 4. Overview of the astroCAMP co-design formulation.
On the left side of Fig. 4, we see how each configuration $x$ is evaluated on a fixed workload $w$ (e.g. an SKA-Low or SKA-Mid benchmark) and is represented as a point in the design space of algorithmic, hardware platforms, and system parameters.
Metrics. For the DSE with astroCAMP framework, shown in Fig. 4, we consider the following core metrics detailed in Table 2:
$T_{c}(x),\;E_{c}(x),\;C_{c}(x),\;\Theta(x),\;C_{TCO},\;Q(x).$
Here, $T_{c}$ is time-to-solution, $E_{c}$ is Energy-to-solution, $C_{c}$ is Carbon-to-solution, $\Theta$ is throughput (e.g., visibilities/s or pixels/Joule), $C_{TCO}$ is the total cost of ownership, and $Q(x)$ is a quality tuple (e.g., dirty-image RMS noise, PSNR). A high-fidelity reference configuration defines $Q_{\mathrm{ref}}$. We measure scalarised quality loss as:
$\Delta Q(x)=d\!\left(Q_{\mathrm{ref}},\,Q(x)\right),$
where $d$ is a weighted distance in quality space.
Multi-objective optimisation problem. We formalise co-design as a multi-objective optimisation problem subject to scientific and operational constraints.
(1) $\min_{x\in\mathcal{X}}\quad\bigl(T_{c}(x),\;E_{c}(x),\;C_{c}(x),\;\Delta Q(x)\bigr)$
Scientific quality constraint:
$\Delta Q(x)\leq\Delta Q_{\max}.$
SDP operational constraints:
(2) $P_{\mathrm{avg}}(x)$ $\leq P_{\max}$ $\text{(site power cap, e.g.\ 2 MW)},$
(3) $\Theta(x)$ $\geq\Theta_{\min}$ $\text{(minimum throughput / survey cadence)},$
(4) $x$ $\in\mathcal{X}_{\mathrm{valid}}$ $\text{(resource bounds, feasibility)}.$
Constraints couple algorithmic, hardware, and workflow choices to architectural, sustainability, and economic layers through energy, power, and carbon metrics. The formulation is modular: new algorithmic parameters extend $a$, new hardware platforms extend $h$, and new workflow or runtime options extend $s$. Additional metrics can be added to the objective or constraints without changing the optimisation structure. In this way, astroCAMP defines a stable design space schema that still accommodates future algorithmic and architectural innovation.
Because the objectives in Eq. (1) conflict, no single optimal configuration exists. We therefore compute the Pareto frontier
$\mathcal{P}=\Bigl\{x\in\mathcal{X}_{\mathrm{valid}}\;\Big|\;x\text{ is not dominated across }(T_{c},E_{c},C_{c},\Delta Q)\Bigr\},$
where “not dominated” means that there is no $x^{\prime}\in\mathcal{X}_{\mathrm{valid}}$ such that $T_{c}(x^{\prime})\leq T_{c}(x),E_{c}(x^{\prime})\leq E_{c}(x),C_{c}(x^{\prime})\leq C_{c}(x),\Delta Q(x^{\prime})\leq\Delta Q(x),$ with at least one inequality strict.
Each $x\in\mathcal{P}$ represents a distinct, scientifically valid, and operationally feasible trade-off (e.g., “fast but power-hungry”, “high precision but slower”, “energy-efficient but low throughput”). The construction of $\mathcal{P}$ is done via structured parameter sweeps at the software and/or the hardware architecture level. Evaluation of co-design metrics can be performed using simulation tools (Sarkar and Hao, 2023), model-based design (Suriano et al., 2017) or surrogate models. Automated design space exploration (Zaourar et al., 2023) can rely on Bayesian optimization (Fu et al., 2025), or evolutionary search to analyze how both new (BIPP, WS-Snapshot) and legacy algorithms (IDG) populate the multi-metric design space.
5. astroCAMP Benchmark Suite and Datasets
The benchmaks and datasets are designed to be large enough to stress memory, I/O, and communication layers while remaining manageable for design-space exploration experiments. We provide open, standardised benchmark cases, representative of radio-interferometric imaging workloads. Each case includes:
  • Synthetic and pathfinder datasets derived from realistic telescope configurations (e.g. SKA-Low), generated with OSKAR;
  • Reference outputs from validated CPU implementations (e.g. WSClean, IDG);
  • Configuration files specifying baseline distribution, frequency range, integration time, and field of view.
The datasets were generated using OSKAR (Mort et al., 2010) (version OSKAR-2.11.2-dev 2025-07-17 ddb65ed) using SKA-Low 512 stations full configuration. Sixteen datasets were produced varying the number of timesteps (or time samples) over 1, 8, 64, 128, and 256, and the number of channels over 1, 8, 64, 128, and 256. Each timestep is integrated over 10 seconds. Phase center was set to 25.0 and -30.0 degrees in right ascension and declination, respectively. Start frequency was set to 151 MHz with increments of 1 MHz. Sources were selected from the GLEAM catalog (Hurley-Walker et al., 2017) to cover a field of view of 40 degrees. Data volumes range from megabytes to terabytes, enabling tests from workstation to cluster scale. Each dataset includes:
  • Raw power and timing logs at the kernel and pipeline level, together with energy metrics.
  • Ground-truth dirty image to compare algorithmic quality metrics.
  • Configuration templates for CPU and GPU runs.
6. Benchmark Evaluation and Results
This section benchmarks SKA-Low workloads and datasets to demonstrate the practical impact of our co-design methodology. We report results from a subset of metrics (performance, energy efficiency, sustainability, and economy), then analyse heterogeneous WSClean+IDG execution, CPU-only strong scaling, location-dependent carbon/cost efficiency for identical workloads, and PREESM-based design-space exploration for performance–energy–utilization trade-offs toward Pareto-optimal SKA-scale operating regions.
6.1. Experimental Setup
Benchmarks were executed on a Lenovo ThinkSystem SR675 V3 node of the Green500 KUMA cluster (SCITAS, ), with dual AMD EPYC 9334 CPUs, 371 GB RAM, 6.4 TB NVMe, and four NVIDIA H100 GPUs (94 GB). All runs used exclusive node access and allocated 16 CPU cores, one H100 GPU, and one quarter of RAM (92.2 GB; Slurm limit 5900 MB/core).
The server cost was obtained by configuring a system with comparable specifications, resulting in $29\,875.39\text{\,}\mathrm{U}\mathrm{S}\mathrm{D}$ (Lenovo, 2024). Embodied hardware emissions were estimated by combining the emissions reported by NVIDIA for the H100 GPU ( (NVIDIA, 2023)) with the emissions of the remaining system components modeled using Boavizta’s Datavizta tool (Boavizta, 2024), yielding a total embodied emission of . The latter assumes a Genoa-class CPU architecture, Samsung memory modules, and a Micron SSD. We assume that the CAPEX is amortized over a six-year lifetime.
Figure 5. Profiling of WSClean + IDG on GPU: 256 timesteps, 256 channels, full SKA-Low, for a 16k $\times$ 16k pixels image.
The data were processed with WSClean (Offringa et al., 2014) in IDG (Van der Tol et al., 2018a) GPU mode. For each of 16 datasets, we fixed FoV and varied image size ($4096^{2}$, $8192^{2}$, $16\,384^{2}$, $32\,768^{2}$; pixel scales 17.578, 8.789, 4.394, 2.197 arcsec). For each of the $25\times 4=100$ runs, we executed: warmup, 120 s pause, monitored run, 120 s pause. Warmup compiles CUDA kernels; the monitored run reuses compiled kernels. The pre-run pause allows hardware to settle. Power measurements were recorded using node-level PDUs, which provided a 1 s power average sampled every 5 s. The PDU measurements were cross-validated with software monitors (RAPL, NVML), using the PMT library wrapper (Corda et al., 2022a). Fig. 5 exemplifies CPU utilization, memory footprint, I/O intensity, and CPU/GPU/system power (as deltas from idle). CPU activity is mostly confined to one core (100 %) with short multi-core bursts, indicating poor CPU-side scaling. GPU power remains well below H100 TDP and streaming multiprocessors (SM) utilization appears in short bursts, indicating that work supply is intermittent and/or constrained by memory-access effects.
6.2. Heterogeneous WSClean+IDG Deployment
Figure 6. Energy hierarchy for image size $32768^{2}$ across all $(n_{\mathrm{times}},n_{\mathrm{chans}})$ configurations. Stacked bars show component-level energy contributions (CPU, GPU, IDG host/device), system overhead, and the corresponding PDU-level measurement. The system total (orange dashed line) and PDU total (blue dashed line) highlight the rack-level overhead above node-level measurements. Execution times (right axis, log scale) include wall time, IDG host time, and IDG device time.
Figure 7. Energy-throughput SKA-Low (WA). Each panel shows a stacked energy breakdown (dynamic + static, in Wh; left axis) alongside throughput (Mvis/h; right axis, log scale), across combinations of timesteps and channels.
Figure 8. Memory-efficiency and footprint trends. Lines report memory efficiency for GPU vs CPU execution of IDG.
Figure 9. Largest-image saturation signals grouped by channels $c$. Left: summed explicit IDG sub-kernel times (gridder, sub-FFT, w-tiling) versus end-to-end WSClean wall time for each $(t,c)$ workload. Right: throughput with active-window GPU utilization and memory-footprint statistics (averaged only over samples with active GPU work).
Figure 10. GPU-vs-CPU roofline comparison for WSClean stacking ($t=256$, $c=256$, $16384^{2}$). The figure contrasts arithmetic intensity, achieved throughput, and bandwidth context for CPU-only execution and the GPU-enabled kernels. Roofline markers shape identifies the workload and color the parallelism.
Figure 11. CPU-only roofline context for WSClean stacking ($t=256$, $c=256$, $16384^{2}$) across $p=\{1,16,32,64\}$: arithmetic intensity, achieved throughput, achieved memory bandwidth, and machine ceilings.
WSClean+IDG is a heterogeneous pipeline: the CPU performs batching, metadata preparation, and scheduling, while the GPU executes IDG gridding and interpolation. Throughput is high only when the CPU feeds work fast enough to keep the GPU busy. Fig. 6 shows a clear hierarchy for $32768^{2}$ across $(n_{\mathrm{times}},n_{\mathrm{chans}})$: GPU and IDG-device energy dominate (about 60–90%), while CPU, IDG-host, and system overheads are smaller and comparatively flat. GPU energy rises from about 5–10 kJ to about 165 kJ at t256t256c256c256, whereas CPU/host-side IDG energy remains nearly constant, confirming accelerator-driven scaling. Execution time and energy grow together over nearly four orders of magnitude, and channel-heavy workloads deliver similar throughput with lower wall time and total energy than time-heavy cases.
[…]
We analyze with Fig. 7 how energy and throughput vary across the astroCAMP dataset configurations. As visibility dimensions increase, throughput improves initially but rapidly saturates, while wall time and total energy grow super-linearly. This trend reflects increasing data-movement and orchestration overheads for larger working sets, rather than a change in the fundamental computational bottleneck. Consequently, naïvely scaling imaging parameters increases operational cost and carbon footprint without delivering commensurate scientific throughput, underscoring the need for efficiency-aware configuration selection for SKA-scale deployments. Darshan profiling shows that I/O contributes only a small share of total runtime ($\sim 1\%$) for this workload, but is composed of many short operations. For the largest case, we also observe many sub-megabyte reads and a high fraction of misaligned requests, indicating that access granularity and request alignment are relevant optimization targets even when peak filesystem bandwidth is not saturated. Data-movement efficiency metric (Bytes/J) quantifies the energetic cost of memory traffic. Complementing this, Fig. 8 provides an analytical decomposition of transferred workload size and shows that scaling is increasingly input-dominated at larger $(t,c)$, which motivates optimization of ingest, staging, reuse, and data-format/compression choices.
Fig. 8 shows payload-footprint and memory-efficiency trends at $32768^{2}$: increasing channels mainly grows input payload, while the output term remains fixed per image size (0.0625 GiB for $4096^{2}$, 0.25 GiB for $8192^{2}$, 1.0 GiB for $16384^{2}$, and 4.0 GiB for $32768^{2}$). The GB/J curves indicate that larger spectral workloads do not translate into proportional memory-efficiency gains; the device path is generally above host, but both are sublinear versus footprint growth. The co-design implication is to prioritize channel-aware batching, fewer redundant visibility passes, lower orchestration overhead, and scheduling that keeps efficient GPU phases dominant for longer.
The analysis in Fig. 9 shows that increasing $c$ raises explicit IDG kernel work, but end-to-end WSClean wall time remains much larger than summed sub-kernel times, so application-level gains are not proportional. Throughput flattens before active-window SM utilization saturates and while active memory utilization stays modest, indicating a workflow-level bottleneck. This matches Nsight Compute: some kernels are efficient while active, but those states are not sustained across the full workflow timeline, so gains are more likely from batching, overlap, synchronization reduction, and staged-data reuse tha n from kernel-only tuning.
The roofline study in Figs. 10 and 11 adds model-guided context for CPU-only versus GPU-enabled execution. The CPU-only path remains to the right of the DRAM ridge point for this workload while achieving only a modest fraction of peak compute, and the GPU-enabled path shows efficient kernels only in intermittent windows. Together, the end-to-end limits are better explained by orchestration, synchronization, and host-device pipeline effects than by a single DRAM-roof constraint.
6.3. Strong CPU-only Scaling Limitations
[…]
Figure 12. WSClean+IDG on 1–64 CPU threads. Bars show absolute runtime, solid line shows speedup relative to 1 thread, and dotted line indicates ideal linear scaling.
Figure 13. Strong-scaling speedup and parallel efficiency for CPU-only WSClean vs GPU accelerated gridding.
Table 3. Execution times for $t=256$, $c=256$, $16384^{2}$ with IDG on CPU or GPU across CPU core counts.
Solution 1c 16c 32c 64c
WSClean + IDG CPU 3:13:10 1:08:28 1:08:10 1:08:06
WSClean + IDG GPU 1:09:27 1:09:57 1:09:14 1:08:45
Figure 14. Darshan summary for a CPU-only run (16 cores), showing I/O runtime share and operation counts.
Table 3 confirms this behavior on the flagship case: CPU-only runtime improves from one to 16 cores, then nearly plateaus, while GPU-enabled wall time remains nearly constant across CPU core counts because the dominant parallel work is already offloaded.
Although bandwidth-dominated kernels can strong-scale in principle; here, losses are dominated by serial regions and orchestration overheads in the full application path. Fig. 13 quantifies this collapse via speedup and parallel efficiency and indicates substantial synchronization overhead from four or more cores. Flamegraph traces also expose libcasa_tables and libcasa_casa in effectively single-threaded WSClean sections, creating a serial metadata/I-O lane that increases waiting time and weakens scaling gains from extra CPU cores. In our runs we use MSv2; moving to MSv4-capable I/O frameworks should reduce this serialization pressure (e.g., better concurrent access and request aggregation), improving pipeline-level scaling in addition to raw I/O efficiency.
Investigating further, we identify that the spin and synchronization overhead grows quickly with thread count, consistent with the observed speedup collapse. Fig. 14 shows that I/O accounts for a small fraction of total wall time ($\sim 1\%$), but with many short and frequently misaligned requests. For the largest case, we observe 99.66% misaligned file requests and 280,024 read operations below 1 MB (262,144 targeting the same data file), indicating that request aggregation and alignment remain relevant even when peak filesystem bandwidth is not saturated.
6.4. Location-Dependent Efficiency for SDPs
Figure 15. Carbon (top) and cost breakdown (bottom) with efficiency metrics for image size (32768²) across all configurations, comparing SKA-Low (WA) and SKA-Mid (SA).
In Fig. 15, we compare carbon and cost for the largest image size at the SKA SDP locations. Throughput is identical, but efficiency diverges with local grid intensity and electricity price. Operational emissions dominate carbon (90-95%); WA attains higher carbon efficiency (Mvis/kgCO2) because grid intensity is lower (0.321 vs. 0.672 kg CO2/kWh). Cost efficiency is CAPEX-dominated (89-92%): WA’s higher electricity price (0.27 vs. 0.19 \$/kWh in SA) yields lower Mvis/\$. Increasing timesteps/channels amortizes embodied carbon and CAPEX, so larger batches improve efficiency.
6.5. Design Space Exploration with PREESM
PREESM (Suriano et al., 2017) is a prime example of a DSE framework to study its extendability. The diversity of metrics, specialized for a scientific field, calls for automated exploration support. The current flow supports many system and hardware platform-level metrics, and more can be computed from those metrics based on user needs, using an existing model for fast evaluation.
Taking a step back to look at metric implementations, algorithmic metrics require the complete execution of the pipeline, with potentially long execution, incompatible with DSE. Future flows should model parameter influence on expensive metrics to avoid systematic evaluation (Honorat et al., 2022). For example, changing quantization may require full execution, while changing hardware frequency may not.
Sustainability and Economic metrics are two examples of environmental metrics, based on external information. These metrics require interaction with external models or databases, enabling a multi-model co-design flow.
We evaluate some of these metrics with PREESM by implementing the simulator, a subset of the DDFacet radio-astronomy pipeline that inverts an artificial sky image into visibilities. We run our designs on a KRIA KR260 SoM board with a 4-core, 1.6 GHz ARM Cortex-A53 processor and an Ultrascale+ FPGA clocked at 150 MHz. We measure each design’s resource usage, latency, and energy-to-completion at execution. The energy is obtained by measuring the base consumption of the device at idle and subtracting it from the consumption under load. Fig. 16 summarizes our results. We define the occupancy metric as the weighted average between the percentage of cores used and the utilization of three FPGA resources: CLBs, DSPs, and BRAMs : $avg(CPU,avg(CLB,DSP,BRAM))$. This metric aims to measure the proportion of total available computing resources used. The FPGA-mappable computation is the 2-D FFT, split into two 1-D FFTs. The application is run on a 128x128 input image, resulting in a 128x128-point 2-D FFT, and outputs 9̃8k visibilities.
Using PREESM, we can explore the design space and derive the Pareto front shown in Fig. 16. This enables informed design decisions—such as the degree of parallelism and the mapping of tasks to CPU or FPGA—by explicitly trading off target objectives, here latency and energy consumption.
[…]
6.46.46.66.66.86.877202040406060latency (s)Occupancy (%)
Figure 16. Pareto front generated with PREESM. Square dots represent CPU-FPGA designs; round dots, CPU-only designs.
7. Discussion of Benchmark Evaluation
The benchmark results support a clear result-to-decision chain for SKA imaging co-design. Our evaluation of WSClean+IDG yields two evidence-based conclusions:
  1. (1)
    Pipeline-level scaling limits dominate: CPU strong scaling collapses after low thread counts (Figs. 12, 13), CPU activity is bursty and mostly single-core in critical stages (Fig. 5), and end-to-end wall time remains far above summed kernel time even when active-window GPU utilization is nonzero (Fig. 9). Together with the roofline comparison (Fig. 10), this points to orchestration/synchronization and serial-path effects rather than a single kernel-bound limit.
  2. (2)
    Efficiency gains are configuration- and site-dependent: energy hierarchy and throughput trends (Figs. 6, 7) show accelerator-dominated scaling with persistent static-energy overheads; payload trends indicate input-dominated growth at large $(t,c)$ (Fig. 8); and carbon/cost efficiency diverges by location despite equal throughput (Fig. 15).
Practical takeaway: use astroCAMP in three steps: (i) detect scaling/utilization bottlenecks (Figs. 5, 12, 13, 9), (ii) compare energy/carbon/cost efficiency across $(t,c)$ and sites (Figs. 7, 15, 8), and (iii) select the smallest configuration that meets science-quality tolerances.
7.1. Why We Need HW–SW Co-Design?
[…]
7.2. Open Challenge for the SKA Community
How Much Quality can we Trade off for Efficiency? For each SKA Key Science Programme—Cosmic Dawn/EoR, Galaxy Evolution and Cosmology, Cosmic Magnetism, and Time-Domain Astrophysics—define quantitative application-level quality metrics and tolerances. Examples include PSNR/SSIM, flux-scale accuracy, astrometric precision, polarisation purity, RM recovery error, spectral-line fidelity, time-series accuracy, and transient completeness. Tolerances must be precise enough to guide hardware–software co-design (precision choices, approximations, accelerators) while ensuring validity for: 21 cm EoR power spectra and tomographic cubes, all-sky continuum and deep extragalactic surveys, RM Grid measurements of cosmic magnetism, high-cadence transient and pulsar timing searches.
Why is this challenge needed? SKA documents define high-level goals—e.g. 1 % astrometry and spectral dynamic ranges of $10^{5}$$10^{4}$—but not survey-level tolerances for flux accuracy, completeness, PSF residuals, polarisation purity, or spectral-line fidelity. Without these thresholds, the co-design community cannot determine which approximations (e.g., reduced precision, coarse $w$-stacking) are acceptable. With SKA power and cost envelopes fixed years ahead, the key question is which algorithm/hardware combinations maximize performance and energy efficiency without violating fidelity. This requires cross-layer metrics linking runtime, energy, carbon, utilization, and image quality, plus programme-level tolerances that are still missing (Bacon et al., 2020).
Community engagement plan. astroCAMP follows an open-source, maintainer-guided model, hosted as a public repository (SEAMS Project, 2025) where datasets, metrics, and reference implementations are curated by maintainers and extended via standard contribution workflows. This mirrors established HPC software practices, supporting transparent validation and gradual community convergence on shared quality metrics and tolerances.
8. Conclusions
We presented astroCAMP as an open benchmark and co-design framework for SKA imaging that couples representative datasets with unified runtime, energy, carbon, and cost metrics. The reported results show: (i) end-to-end limits are dominated by orchestration/synchronization and serial-path effects, not only kernel-level ceilings (Figs. 5, 9, 10); (ii) CPU-only scaling is poor for SKA-sized workloads (Figs. 12, 13); and (iii) sustainability/economic outcomes depend on workload configuration and site context even at equal throughput (Figs. 7, 15). These findings translate into a practical strategy: prioritize pipeline utilization and data movement efficiency, then optimize architecture and scheduling under explicit quality tolerances. We conclude with a community call to: (1) establish science-driven imaging-quality tolerances, (2) expand benchmark datasets, (3) develop reference submissions for current and emerging imaging pipelines, and (4) build an open submission and validation process for the SKA era.
9. Acknowledgments
This research was funded, in whole or in part, by the Agence National de la Recherche grant agreements ANR-23-CE46-0010 and ANR-22-EXNU-0004, and the Swiss National Science Foundation grant no. 200021E_220194: “Sustainable and Energy Aware Methods for SKA (SEAMS)”. A CC BY license is applied to the AAM resulting from this submission, in accordance with the open access conditions of the grant. The authors thank the EPFL EcoCloud center, in particular Dr. Xavier Ouvrard, for providing access to its infrastructure for monitoring energy consumption in servers.